OpenStudy (anonymous):

Sketch and find the area of the region bounded by the given curves. y = x^2 + 1, y = 3x - 1

OpenStudy (luigi0210):

Graph the equations

OpenStudy (anonymous):

how?

OpenStudy (luigi0210):



OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the area between those two lines is the bounded region right?

OpenStudy (anonymous):

by the way is the graph in the first quadrant?

OpenStudy (luigi0210):

Yes, and yes it is in quad one

OpenStudy (anonymous):

okay so what do i do next? after I graph it? do I integrate?

OpenStudy (amistre64):

yes; integrate the difference of the functions along the interval in which they are bounded

OpenStudy (anonymous):

sooo i'll be integrating x^2 + 1 and 3x - 1?

OpenStudy (amistre64):

determine when: f(x) = g(x) to get the interval of [a,b] then integrate f-g, from a to b

OpenStudy (amistre64):

if you come up with a negative, drop the sign :)

OpenStudy (luigi0210):

Yup, and make sure you are subtracting the top from bottom

OpenStudy (amistre64):

top from bottom is superfluous

OpenStudy (amistre64):

5 - 3 = 2 3 - 5 = -2 your result should be positive, so the only difference is the sign

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i'm integrating x^2 + 1 and 3x - 1? then subtracting top with the bottom?

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@RadEn

OpenStudy (dumbcow):

you are integrating the difference of 2 functions

OpenStudy (anonymous):

ohh so i'll subtract then integrate the difference of them?